Suppose the monthly demand for a product is given by 10 100 x p e   where p is the price per unit when x units are sold

(i) What is the total revenue for the commodity?

(ii) What would be the total revenue when30 units were demanded and supplied?


1
Expert's answer
2021-09-13T00:29:05-0400

Solution:

Since, given demand function is unclear, we assume, "P(x)=100x-10"

(i) "R(x)=x.P(x)=x(100x-10)=100x^2-10x"

(ii) Put "x=30"

"R(30)=100(30)^2-10
\\=100(900)-10
\\=90000-10
\\=89990"